OpenStudy (anonymous):

the shorter leg of a 30-60-90deqree triangle is 6. what is the lenght of the other leg?

OpenStudy (anonymous):

HELP ME PLEASE. :)

OpenStudy (anonymous):

here sin 30 deg = 6/B or 1/2 = 6/B or B = 12 so length of other leg is 12

OpenStudy (anonymous):

So the answer is12?